OpenStudy (anonymous):

If 'x' is the level of the killed hero, the amount of experience gained is: f(x) = f(x-1) + 20x; f(0) = 220; What is the experienced gained if the killed hero is level 5?

OpenStudy (anonymous):

We can do this recursively. Start by finding what f(1) is. Since you know what f(0) is, you can calculate f(1) by plugging in 1 for x and substituting f(0).

OpenStudy (anonymous):

\[ \begin{split} f(x) &= f(x-1)+20x\\ &= (f(x-2)+20x)+20x\\ &= f(x-2)+2\cdot 20x \\ &=f(x-x) +x\cdot 20x\\ &=f(0) + 20x^2\\ f(x) &= 20x^2+220 \end{split} \]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Hold on, what I did was not quite correct though.

OpenStudy (anonymous):

Since I should have put \(f(x-1)=f(x-2)+20(x-1)\).

OpenStudy (anonymous):

So \[ \begin{split} f(x) &= f(x-n) + 20x\sum_{k=1}^{n}x+1-k\\ &=f(x-n) +20x\left(n(x+1)+\frac{n(n+1)}{2}\right) \end{split} \]

OpenStudy (anonymous):

This is a bit trickier, so it may be easier to just it manually.\[ f(1) = f(0) + 20(1) = 220+20 = 240\\ f(2) = f(1) + 20(2) = 240+40 = 280\\ \dots \]

OpenStudy (anonymous):

ok thank you for all your help i appreciate it